5x + 8 + 3x - x + 5 = 6x - 3
Answer by mathslover(132)

The given equation is :
5x + 8 + 3x - x + 5 = 6x - 3
the way to solve such equations is first to group all the variable terms (here x) at one side and the constants on the other.
So we have :
5x + 3x - x - 6x = -8 - 5 - 3
x = -16